DCC and Trigger Readout of SDC Straw Tube Tracker

• Introduction

• Overview of Proposed DCC

• Functionality of Tracking Trigger & Readout • Proposed Layout of Trigger Readout

Introduction

Conceptual baseline design for DCC system now exists

• ASD’s reside on separate cards

• Signals from each module routed to μcrate

• Data link from each μcrate to Crate Interface Card (CIC)

• 8 μcrates / CIC and 12 CIC’s / SDC DAQ Crate

• 2 DAQ Interfaces / DAQ Crate

Trigger data readout has remained open question

• Considered using DCC as datapath – incompatible with L1 trigger data readout – difficult to integrate into data-driven design

• Assumed independent data paths for L1 & L2 through CIC

L1 and L2 Tracking Triggers

Assumed each axial module instrumented with Mean Timers

• Will give at least two Pt thresholds

• 60 MHz, synchronous with beam crossings

• Spatial resolution of ∼ 1 wire diameter

For L1, only require coarse information:

• Baseline requires ∼ 64 azimuthal segments

• Acheived by sending ”OR” of Pt bits over whole module • Gives Δϕ ∼ 2π/200 for outer layer

For L2, require most resolution:

• Location of segment to with a wire

• ∼ 27 possible locations

• Sent out at L1 trigger rate

Implementation of Trigger Readout

Send trigger signals from each μcrate board to interface card

• ∼ 27 signals to chip on interface board

• latch trigger data at each crossing

• form “OR” and generate 2 bits of L1 data

• Store bits for L1 decision

• Transmit all data for L2 trigger

Requires:

• Dedicated 120 Mbit channel from ∼ 700 modules

• 10 Mbit/s data path (Cu?) for L2 data

Data sent to CIC where it is multiplexed

• ⇒ 96 1 Gbit/s fibre-optic L1 inputs

• ⇒ 96 80 Mbit/s L2 inputs